general solution of tan5A=cot2A?

Grade:11

2 Answers

Sunil Raikwar
9 years ago
tan5A= cot2A
tan5A= tan(pi/2-2A)
general solution is
5A=npi+(pi/2-2A)
7A=(2n+1)pi/2
Thanks & Regards
Sunil Raikwar
askIITians faculty
Aman Kumar Jyotishi
6 years ago
Here is the answer for the question,I hope the one who sees this would find it usefultan 5A = tan(90-A)5A = 90 - A90 = 6ATherefore, A = 15